# Adversarial Robustness in Parameter-Space Classifiers

### v1

@@ Abstract
Implicit Neural Representations (INRs) have been recently garnering increasing interest in various research fields, mainly due to their ability to represent large, complex data in a compact and continuous manner. Past work further showed that numerous popular downstream tasks can be performed directly in the INR parameter-space. Doing so can substantially reduce the computational resources required to process the represented data in their native domain. A major difficulty in using modern machine-learning approaches, is their high susceptibility to adversarial attacks, which have been shown to greatly limit the reliability and applicability of such methods in a wide range of settings. In this work, we show that parameter-space models trained for classification are inherently robust to adversarial attacks – without the need of any robust training. To support our claims, we develop a novel suite of adversarial attacks targeting parameter-space classifiers, and furthermore analyze practical considerations of attacking parameter-space classifiers. Code for reproducing all experiments and implementation of all proposed methods will be released upon publication.
@@ 1 Introduction
Implicit Neural Representations (INRs) are representations of an arbitrary signal as a neural network predicting signal values under some conditioning. These representations have recently been raising increased interest due to their ability to efficiently and compactly encode high-dimensional data while maintaining high input fidelity, and have further shown to be capable of learning additional intrinsic knowledge latent within the original signal-domain representation (Sitzmann et al., 2020; Wang et al., 2021).
While INRs are inherently designed to fit one specific signal (rather than generalizing across an entire dataset), several recent works have proposed ways to use these representations to achieve a variety of downstream tasks by using some neural meta-network operating directly over the parameter-space of the INR of each signal within the dataset ((Chen & Wang, 2022; Dupont et al., 2022; Lee et al., 2021; Navon et al., 2023b)).
In these settings, each data sample is received in the native signal space, converted to its compressed INR representation by fitting a neural model, and finally the meta-network performs the designated downstream task over the implicit representation encoded by the model’s parameters. The main advantage of such approaches is that they alleviate challenges in tasks involving high-dimensional data (videos, point-clouds, spectrograms), often requiring the incorporation of large, complex and problem-specific neural architectures for efficient data processing, and also potentially demand large computational resources for efficient optimization.
Adversarial attacks are known to be a major concern for the use of neural networks across a variety of tasks. In this work, we explore – to the best of our knowledge, for the first time – the adversarial robustness of parameter-space classifiers. We start by framing the problem: any attack on a parameter-space classifier must be simultaneously optimized to reduce classification performance while also satisfy some constraint incurring the original, signal domain data fidelity. Since no prior studies have addressed this setting, we first formulate a novel set of possible white-box attacks over several data modalities. We then proceed to show that discussed parameter-space classification models are inherently robust to adversarial attacks in comparison to the signal-domain classification problem counterpart, as exemplified using our established attack suites.
Our hypothesis of inherent robustness can be motivated by the following: Firstly, adversarial attacks commonly utilize the highly-nonlinear nature of modern deep classifiers to create perturbations with specific patterns that, while bounded, will be heavily amplified by the classifier to create significant data manipulations in deeper layers ((Ingle & Kulkarni, 2021; Romano et al., 2020)). In this setting of parameter-space classification, INR optimization is performed per-sample, and is embedded onto the classifier’s inference process. This constrains the attacker to optimize perturbations in the signal-domain rather than directly manipulating the classifier’s input. This optimization is disjoint from the downstream objective, and is dedicated to expressing global data features as a whole, rather than entangle specific entries with specific local signal features (Sitzmann et al., 2020). Therefore, this optimization process tends to obfuscate adversarial patterns introduced in signal domain. Secondly, as we later show in Section 4.4, performing adversarial attacks over parameter-space classifiers poses significant computational difficulties for a white-box attacker, as attacking the INR given a perturbed signal requires backpropagation through an internal optimization loop.
In summary, this paper makes the following contributions:
  1. 1.
    We supply evidence of inherent robustness of parameter-space classifiers to adversarial attacks.
  2. 2.
    We propose several novel types of adversarial white-box attacks for parameter-space classifiers.
  3. 3.
    To demonstrate our claims on 3D data, we also formulate a novel adversarial attack for voxel-grid data representations, and demonstrate its performance in both parameter-space and signal-space classification settings.
@@ 2 Related Work
The analysis of trained neural network parameters had been raising increasing interest in recent years, demonstrating applicability in a wide, and at times surprising, range of machine learning tasks. (Eilertsen et al., 2020) was one of the first works considering this idea, performing trained network parameter analysis to identify the optimization procedure hyperparameters. (Unterthiner et al., 2020) showed that a trained CNN classifier model parameters can be used to directly attain a good assessment of classification accuracy, without running the full model’s forward pass. (Dupont et al., 2022) have been the first to demonstrate that neural network parameters, specifically of the SIREN (Sitzmann et al., 2020) models (one of the first INRs to be proposed), could be used as the direct and only input for the training of a meta-learner model for classification, generative, data-imputation and novel-view synthesis tasks. (Peebles et al., 2022) employ generative modeling of deep-weight space to improve parameter update steps in model training.
@@ 3.2 Parameter-Space Classifiers
In contrast to conventional classifiers operating directly on the signal space $\mathcal{X}$, parameter-space classifiers operate on the signal INRs, $R(\mathcal{X})\subset\mathcal{Y}$. We denote a parameter-space classifier as $M_{\psi}$, parameterized by $\psi$; the model receives the representation parameters $\theta=R(x)$ as the input and outputs the predicted class label $\hat{c}=M_{\psi}(\theta)$ as the output.
Even for small neural representation models, the parameter space $\mathcal{Y}$ is often high-dimensional, increasing difficulties related to computational resources and overfitting when training the classification models on it (Navon et al., 2023a, b). Functas (Dupont et al., 2022) cope with this issue of increased parameter-space dimensionality by partitioning the parameter space into weights and biases, $\mathcal{Y}=\mathcal{Y}_{\mathrm{w}}\cup\mathcal{Y}_{\mathrm{b}}$, and sharing a single vector of weights $\theta_{\mathrm{w}}\in\mathcal{Y}_{\mathrm{w}}$ across a collection of represented signals, while allowing to fit individual modulation vectors $v(x)\in\mathbb{R}^{d}$ of a fixed dimension with a learned shared linear decoder $W$ such that each signal has an individual vector of biases, $\theta_{\mathrm{b}}(x)=Wv(x)$. This technique was successfully demonstrated with the SIREN (Sitzmann et al., 2020) models.
In what follows, we continue denoting the representation of $x$ as $\theta$, while taking into account that the actual INR $F$ may have additional parameters shared across multiple signals.
@@ 3.3 Parameter-Space Adversarial Attacks
Note that while the signal is attacked in its native domain, the classifier $M$ accepts its input in the representation domain. The two domains are related by the map $R$ – a function whose pointwise value requires the solution of an optimization problem via a numerical algorithm. While backpropagating through this optimization process is possible (e.g. via automatic differentiation or implicit differentiation (Navon et al., 2020)), as we later show in Section 4, doing so dramatically increases computational resources required for perturbation optimization, thus posing additional practical challenges for a potential attacker, in addition to the inherent difficulty of optimization such an attack that stems from the inference pipeline itself, as motivated in section 1.
@@ 3.3.1 Full PGD
A straightforward approach to solving equation (2) is by performing projected gradient descent (PGD) and backpropagating through both the classification model $M$ and the modulation optimization steps collectively denoted $R(x)$. While backpropagation through $M_{\psi}$ can be performed based on common first-order differentiation methods, a potential attacker must allow gradient flow through the modulation optimization steps performed in the forward pass of $R$. This requires differentiation through those optimization steps, namely via second-order differentiation and optimization. As we further show in Section 4.4, this may substantially increase the amount of computational resource required for the attacker.
@@ 3.3.4 ICOP
Proposed attacks in previous sections focus on perturbations applied in signal domain. In this section we further propose an attack applied in INR domain. The main challenge here is efficiently constraining the adversarial perturbation - while we perturb the neural representation of the signal, we still must constrain the response of every applied perturbation in signal domain. Namely, unlike the formulation in equation 2, here we must impose $F(\delta)\leq B$. The need to query the Functa model $F$ to assess this response raises a difficulty to apply any trivial, immediate projection of a given perturbation $\delta$ onto the feasible set. We therefore propose Imposition of Constraints via Orthogonal Projection (ICOP) - this method relies on two approaches to impose soft constraints over the perturbations - First, after every PGD step we compute a projection loss, measuring constraint violation in $\mathcal{L}_{1}$: $\mathcal{L}_{proj.}=\|\mathcal{L}_{q}(F(\delta))-B\|_{1}$. We then iteratively take small gradient steps $\eta$: $\delta_{proj.}=\delta-\eta\frac{\partial\mathcal{L}_{proj.}}{\partial\delta}$, until constraint satisfaction.
Secondly, to discourage perturbation steps from severely breaking signal-domain constraints, we also found it useful to project the maximized classification loss to be orthogonal to $\frac{\partial\|F(\theta+\delta)-F(\theta)\|_{1}}{\partial\delta}$ (namely, the perturbation-induced deviation in signal-domain) in every perturbation gradient step. We enforce soft-constraints in optimization, and then impose hard constraints over the optimized perturbation to ensure feasibility. This is done by computing $R$ over $F(\delta)$ clamped onto the feasible set.
@@ 3.3.5 Implicit Differentiation
Implicit differentiation is a technique used to differentiate equations where the dependent and independent variables are intertwined, allowing for gradient computation in optimization problems where the input and output cannot be explicitly separated. In our adversarial setting (Equation 2), while $x$ and $M_{\psi}(R(x+\delta))$ are separable, any explicit differentiation method (e.g. computational graph-based gradient computation, as is the case in all previously proposed methods) would be heavily hindered by the nature of $R$. The successive optimization steps required for querying $R$ may vastly inflate the computational graph maintained for the adversarial objective, leading to increased computational costs and potential numerical instability commonly associated with large graphs.
To this end, we also investigate implicit differentiation within our setting. Let $d(x,\delta)$ denote the classification error (as featured in Equation 2), and $\mathscr{L}(\theta^{*},x)$ represent the INR signal approximation error $\|F(\theta^{*})-x\|_{\mathcal{X}}$.
@@
Note Equation 3 circumvents the evaluation of $\frac{\partial\theta^{*}}{\partial x}$ via the chain-rule , and therefore explicit backpropagation through the optimization steps applied when solving Equation 1 is not needed. Therefore, unlike previous methods proposed in this paper (Sections 3.3.1, 3.3.2, 3.3.3 and 3.3.4), the size of the computational graph remains constant with growth in the number of modulation optimization steps. In spite of its merits, we stress that implicit differentiation does not nullify the need of explicit methods – the efficacy of implicit differentiation is heavily reliant on the extent of satisfaction of the underlying optimality conditions. Specifically, in our setting, implicit differentiation is only reliable when the set of parameters $\theta^{*}$ given by $R(x)$ satisfies sufficient signal-domain fidelity with $x$.
@@ 3.4 BVA - Voxel-Grid Space Adversarial Attack
To study adversarial attacks over 3D structures, we must choose how to represent them, so that we can establish appropriate attack perturbation and constraints. Out of the 3 most common conventions for 3D data representations – voxel-grids, mesh-grids and point-clouds – in this work we specifically study adversarial perturbations over voxel-grid data. The reason for this choice is that our experiments show that due to the highly irregular structure of mesh-grids and point-clouds, fitting the dataset-shared set of INR weights ($\theta_{w}$ in notation from section 3.1) yields poor results, preventing achieving implicit representations of adequate signal fidelity.
While adversarial attacks over 3D have been studied in the past for mesh-grids ((Xu et al., 2022)) and point clouds (Liu et al., 2019a), to the best of our knowledge no similar work had so far been done in the context of voxel-grid representations. Since we wish to perform adversarial attacks over INR classifiers of voxel grids, and compare robustness to a corresponding signal-domain classifier, we develop Binary Voxel Attack (BVA) - a novel adversarial attack applied over voxel-grid data.
@@
where $\delta\in\{0,1\}^{H\times W\times D}$ is the optimized binary adversarial perturbation, and $\oplus$ denotes the bit-wise xor operation (performing bit switch over $x$ entries where $\delta$ has logical value of $1$). $\delta$ is chosen to be bounded in $\mathcal{L}_{0}$ norm, namely bounding the number of bits flipped in $x$.
This formulation calls for the incorporation of binary optimization techniques, as trivial binarization of $\delta$ (e.g. by thresholding) is non-differentiable. To optimize $\delta$ as a binary mask we follow the binary optimization technique presented in (Iliadis et al., 2020). This methods maintain two congruent masks - a binary mask $\Phi$ and a real mask $\Phi_{c}$. $\Phi$ is used for forward and backpropagation, where each gradient step applies the gradients calculated for $\Phi$ over $\Phi_{c}$. The latter is then binarized (without gradient tracking) to produce $\Phi$ for the next iteration.
To demonstrate inherent robustness of parameter-space classifiers over voxel-grid data, in section 4.2.1 we first exemplify the efficacy of signal-domain attacks using BVA. Then, in section 4.2.2 we employ BVA in parameter-space attacks and demonstrate increased adversarial robustness. We emphasize that equation 4 is a specific instance of equation 2, and therefore all parameter-space attacks previously proposed in section 3 can be performed on voxel-grids using BVA.
@@ 4 Results
In the following sections we demonstrate the inherent robustness of Parameter-Space classifiers over 3 different datasets - MNIST ((Deng, 2012)), Fashion-MNIST ((Xiao et al., 2017)) and ModelNet10 ((Wu et al., 2015)).
MNIST and Fashion-MNIST are chosen to represent the prevalent and heavily-studied modality of 2D images. While current adversarial robustness and computer-vision literature tends to focus on more challenging and diverse datasets (e.g. CIFAR100, ImageNet), applicability of these datasets in the parameter-space classification setting is currently limited. This is because state-of-the-art parameter-space classification algorithms proposed thus far have yet to achieve classification satisfying classification accuracies on such complex datasets ((Navon et al., 2023a, b; Chen & Wang, 2022)). ModelNet is used to demonstrate our findings over one of the modalities where usage of INRs is naturally useful, due to commonly large dimensionality. Following the approach from (Dupont et al., 2022), for each dataset $\mathcal{X}$ we first optimize a set of modulation vectors $\{v_{x}\forall x\in\mathcal{X}\}$. Then, we train a single MLP classifier over this modulation dataset. Finally, every pre-trained classifier is attacked with each of the five attacks presented in Section 3.3. We emphasize no robust are adversarial training is being done in any stage of this experimentation scheme. Attack formulations and results for 2D data are presented in section 4.1, and for 3D data in section 4.2. In section 4.4 we further demonstrate the added computational costs induced by performing adversarial attacks over deep weight-space classifiers, posing added practical challenges for potential attackers. Full training details are further specified in Section A.1.

### v2

@@ Abstract
Implicit Neural Representations (INRs) have been recently garnering increasing interest in various research fields, mainly due to their ability to represent large, complex data in a compact and continuous manner. Past work further showed that numerous popular downstream tasks can be performed directly in the INR parameter-space. Doing so can substantially reduce the computational resources required to process the represented data in their native domain. A major difficulty in using modern machine-learning approaches, is their high susceptibility to adversarial attacks, which have been shown to greatly limit the reliability and applicability of such methods in a wide range of settings. In this work, we show that parameter-space models trained for classification are inherently robust to adversarial attacks – without the need of any robust training. To support our claims, we develop a novel suite of adversarial attacks targeting parameter-space classifiers, and furthermore analyze practical considerations of attacking parameter-space classifiers.
@@ 1 Introduction
Implicit Neural Representations (INRs) are representations of an arbitrary signal as a neural network predicting signal values under some conditioning. These representations have recently been raising increased interest due to their ability to efficiently and compactly encode high-dimensional data while maintaining high input fidelity, and have further shown to be capable of learning additional intrinsic knowledge latent within the original signal-domain representation Sitzmann et al. (2020); Wang et al. (2021). While INRs are inherently designed to fit one specific signal (rather than generalizing across an entire dataset), several recent works have proposed ways to use these representations to achieve a variety of downstream tasks by using some neural meta-network operating directly over the parameter-space of the INR of each signal within the dataset (Chen & Wang (2022); Dupont et al. (2022); Lee et al. (2021); Navon et al. (2023b)). In these settings, each data sample is received in the native signal space, converted to its compressed INR representation by fitting a neural model, and finally the meta-network performs the designated downstream task over the implicit representation encoded by the model’s parameters. The main advantage of such approaches is that they alleviate challenges in tasks involving high-dimensional data (videos, point-clouds, spectrograms), often requiring the incorporation of large, complex and problem-specific neural architectures for efficient data processing, and also potentially demand large computational resources for efficient optimization. Adversarial attacks are known to be a major concern for the use of neural networks across a variety of tasks. In this work, we explore – to the best of our knowledge, for the first time – the adversarial robustness of parameter-space classifiers. We start by framing the problem: any attack on a parameter-space classifier must be simultaneously optimized to reduce classification performance while also satisfy some constraint incurring the original, signal domain data fidelity. Since no prior studies have addressed this setting, we first formulate a novel set of possible white-box attacks over several data modalities. We then proceed to show that discussed parameter-space classification models are inherently robust to adversarial attacks in comparison to the signal-domain classification problem counterpart, as exemplified using our established attack suites. Our hypothesis of inherent robustness can be motivated by the following: Firstly, adversarial attacks commonly utilize the highly-nonlinear nature of modern deep classifiers to create perturbations with specific patterns that, while bounded, will be heavily amplified by the classifier to create significant data manipulations in deeper layers (Ingle & Kulkarni (2021); Romano et al. (2020)). In this setting of parameter-space classification, INR optimization is performed per-sample, and is embedded onto the classifier’s inference process. This constrains the attacker to optimize perturbations in the signal-domain rather than directly manipulating the classifier’s input. This optimization is disjoint from the downstream objective, and is dedicated to expressing global data features as a whole, rather than entangle specific entries with specific local signal features Sitzmann et al. (2020). Therefore, this optimization process tends to obfuscate adversarial patterns introduced in signal domain. Secondly, performing adversarial attacks over parameter-space classifiers poses significant computational difficulties for a white-box attacker, as attacking the INR given a perturbed signal requires backpropagation through an internal optimization loop. In summary, this paper makes the following contributions – First, we supply evidence of inherent robustness of parameter-space classifiers to adversarial attacks. Second, we propose several novel types of adversarial white-box attacks for parameter-space classifiers. Lastly, to demonstrate our claims on 3D data, we also formulate a novel adversarial attack for voxel-grid data representations.
@@ 2 Related Work
The analysis of trained neural network parameters had been raising increasing interest in recent years, demonstrating applicability in a wide, and at times surprising, range of machine learning tasks. Eilertsen et al. (2020) was one of the first works considering this idea, performing trained network parameter analysis to identify the optimization procedure hyperparameters. Unterthiner et al. (2020) showed that a trained CNN classifier model parameters can be used to directly attain a good assessment of classification accuracy, without running the full model’s forward pass. Dupont et al. (2022) have been the first to demonstrate that neural network parameters, specifically of the SIREN Sitzmann et al. (2020) models (one of the first INRs to be proposed), could be used as the direct and only input for the training of a meta-learner model for classification, generative, data-imputation and novel-view synthesis tasks. Peebles et al. (2022) employ generative modeling of deep-weight space to improve parameter update steps in model training.
@@ 3.2 Parameter-Space Classifiers
In contrast to conventional classifiers operating directly on the signal space $\mathcal{X}$, parameter-space classifiers operate on the signal INRs, $R(\mathcal{X})\subset\mathcal{Y}$. We denote a parameter-space classifier as $M_{\psi}$, parameterized by $\psi$; the model receives the representation parameters $\theta=R(x)$ as the input and outputs the predicted class label $\hat{c}=M_{\psi}(\theta)$ as the output. Even for small neural representation models, the parameter space $\mathcal{Y}$ is often high-dimensional, increasing difficulties related to computational resources and overfitting when training the classification models on it (Navon et al., 2023a; b). Functas (Dupont et al., 2022) cope with this issue of increased parameter-space dimensionality by partitioning the parameter space into weights and biases, $\mathcal{Y}=\mathcal{Y}_{\mathrm{w}}\cup\mathcal{Y}_{\mathrm{b}}$, and sharing a single vector of weights $\theta_{\mathrm{w}}\in\mathcal{Y}_{\mathrm{w}}$ across a collection of represented signals, while allowing to fit individual modulation vectors $v(x)\in\mathbb{R}^{d}$ of a fixed dimension with a learned shared linear decoder $W$ such that each signal has an individual vector of biases, $\theta_{\mathrm{b}}(x)=Wv(x)$. This technique was successfully demonstrated with the SIREN (Sitzmann et al. (2020)) models. In what follows, we continue denoting the representation of $x$ as $\theta$, while taking into account that the actual INR $F$ may have additional parameters shared across multiple signals.
@@ 3.3 Parameter-Space Adversarial Attacks
Note that while the signal is attacked in its native domain, the classifier $M$ accepts its input in the representation domain. The two domains are related by the map $R$ – a function whose pointwise value requires the solution of an optimization problem via a numerical algorithm. While backpropagating through this optimization process is possible (e.g. via automatic differentiation or implicit differentiation Navon et al. (2020)), as we later show in Section 4, doing so dramatically increases computational resources required for perturbation optimization, thus posing additional practical challenges for a potential attacker, in addition to the inherent difficulty of optimization such an attack that stems from the inference pipeline itself, as motivated in section 1.
@@ 3.3.1 Full PGD
A straightforward approach to solving equation (2) is by performing projected gradient descent (PGD) and backpropagating through both the classification model $M$ and the modulation optimization steps collectively denoted $R(x)$. While backpropagation through $M_{\psi}$ can be performed based on common first-order differentiation methods, a potential attacker must allow gradient flow through the modulation optimization steps performed in the forward pass of $R$. This requires differentiation through those optimization steps, namely via second-order differentiation and optimization. As we further show in appendix C, this may substantially increase the amount of computational resource required for the attacker.
@@ 3.3.4 ICOP
Proposed attacks in previous sections focus on perturbations applied in signal domain. In this section we further propose an attack applied in INR domain. The main challenge here is efficiently constraining the adversarial perturbation - while we perturb the neural representation of the signal, we still must constrain the response of every applied perturbation in signal domain. Namely, unlike the formulation in equation 2, here we must impose $F(\delta)\leq B$. The need to query the Functa model $F$ to assess this response raises a difficulty to apply any trivial, immediate projection of a given perturbation $\delta$ onto the feasible set. We therefore propose Imposition of Constraints via Orthogonal Projection (ICOP) - after every PGD step we compute a projection loss, measuring constraint violation in $\mathcal{L}_{1}$: $\mathcal{L}_{proj.}=\|\mathcal{L}_{q}(F(\delta))-B\|_{1}$. We then iteratively take small gradient steps $\eta$: $\delta_{proj.}=\delta-\eta\frac{\partial\mathcal{L}_{proj.}}{\partial\delta}$, until constraint satisfaction. To discourage perturbation steps from severely breaking signal-domain constraints, we also found it useful to project the maximized classification loss to be orthogonal to the perturbation-induced deviation in signal-domain $\frac{\partial\|F(\theta+\delta)-F(\theta)\|_{1}}{\partial\delta}$ in every perturbation gradient step.
@@ 3.3.5 Implicit Differentiation
In our adversarial setting (eq. 2), while $x$ and $M_{\psi}(R(x+\delta))$ are separable, any explicit differentiation method (e.g. computational graph-based gradient computation, as is the case in all previously proposed methods) would be heavily hindered by the nature of $R$. The successive optimization steps required for querying $R$ may vastly inflate the computational graph maintained for the adversarial objective, leading to increased computational costs and potential numerical instability commonly associated with large graphs. To this end, we also investigate implicit differentiation within our setting. Let $d(x,\delta)$ denote the classification error (as featured in eq. 2), and $\mathscr{L}(\theta^{*},x)$ represent the INR signal approximation error $\|F(\theta^{*})-x\|_{\mathcal{X}}$.
@@
Note that unlike previous methods proposed in this paper (sections 3.3.1, 3.3.2, 3.3.3 and 3.3.4), eq. 3 circumvents the evaluation of $\frac{\partial\theta^{*}}{\partial x}$ via the chain-rule , and therefore explicit backpropagation through the optimization steps applied when solving eq. 1 is not needed. In spite of its merits, we stress that implicit differentiation does not nullify the need of explicit methods – the efficacy of implicit differentiation is heavily reliant on the extent of satisfaction of the underlying optimality conditions. Under our formulation – when the set of parameters $\theta^{*}$ given by $R(x)$ satisfies sufficient signal-domain fidelity with $x$.
@@ 3.4 BVA - Voxel-Grid Space Adversarial Attack
Out of the 3 most common conventions for 3D data representations – voxel-grids, mesh-grids and point-clouds – we find that due to the highly irregular structure of mesh-grids and point-clouds, fitting the dataset-shared set of INR weights ($\theta_{w}$ in notation from section 3.1) yields poor results, preventing achieving implicit representations of adequate signal fidelity. Therefore, in this work we focus on voxel-grid representations of 3D data.
While adversarial attacks over 3D have been studied in the past for mesh-grids (Xu et al. (2022)) and point clouds Liu et al. (2019a), to the best of our knowledge no similar work had so far been done in the context of voxel-grid representations. Since we wish to perform adversarial attacks over INR classifiers of voxel grids, and compare robustness to a corresponding signal-domain classifier, we develop Binary Voxel Attack (BVA) - a novel adversarial attack applied over voxel-grid data.
@@
where $\delta\in\{0,1\}^{H\times W\times D}$ is the optimized binary adversarial perturbation, and $\oplus$ denotes the bit-wise xor operation. $\delta$ is chosen to be bounded in $\mathcal{L}_{0}$ norm, namely bounding the number of bits flipped in $x$. To perform the required binary optimization of $\delta$ we follow Iliadis et al. (2020), elaborated in section A.3.
@@ 4 Results
In the following sections we experiment with 3 different datasets - MNIST (Deng (2012)), Fashion-MNIST (Xiao et al. (2017)) and ModelNet10 (Wu et al. (2015)).
Following the approach from Dupont et al. (2022), for each dataset $\mathcal{X}$ we first optimize a set of modulation vectors $\{v_{x}\forall x\in\mathcal{X}\}$. Then, we train a single MLP classifier over this modulation dataset. Finally, every pre-trained classifier is attacked with each of the five attacks presented in section 3.3. We emphasize no robust are adversarial training is being done in any stage of this experimentation scheme. Attack formulations and results for 2D data are presented in section 4.1, and for 3D data in section 4.2. In section C we further demonstrate the added computational costs induced by performing adversarial attacks over deep weight-space classifiers, posing added practical challenges for potential attackers. Full training details are further specified in section A.1.